OpenStudy (anonymous):

9+4(2x-1)+8 use distributive property and simplify the expression A 8x+13<<<< B 6x+20 C 2x+20 D 8x+16 @texaschic101

OpenStudy (texaschic101):

9 + 4(2x - 1) + 8 9 + 8x - 4 + 8 8x + 13...you got it

OpenStudy (anonymous):

Oh cool thanks

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

Combine like terms to simplify the expression 5+7+15xy^2+9y^3+10x-13xy^2-3x-8y^3 I think its y^6+2xy^2+7x+12

OpenStudy (anonymous):

Y^3 not y^6

OpenStudy (texaschic101):

5 + 7 + 15xy^2 + 9y^3 + 10x - 13xy^2 - 3x - 8y^3 combine like terms 5 + 7 = 12 15xy^2 - 13xy^2 = 2xy^2 9y^3 - 8y^3 = y^3 10x - 3x = 7x put them together y^3 + 2xy^2 + 7x + 12 --- you got it again...way to go :)

OpenStudy (anonymous):

Wow I also need to see if this is a true statement 18-2m< or equal to 7+3m when m=2

OpenStudy (anonymous):

I think it is

OpenStudy (texaschic101):

18 - 2m <= 7 + 3m (when m = 2) sub in 2 for m 18 - 2(2) <= 7 + 3(2) 18 - 4 <= 7 + 6 14 <= 13 nope...not true....because 14 is not less then or equal to 13

OpenStudy (anonymous):

Oh okay thanks

OpenStudy (texaschic101):

sure thing :)